OpenStudy (samigupta8):

Given vectors a=1/2i +√3/2j And b=√3/2j+1/2k And c=i+2j+3k .The volume of parallelepiped with three coterminous edges as u=(a•a)a +(b•b)b +(c•c)c v=(a•b)a +(b•b)b +(b•c)c w=(a•c)a +(b•c)b +(c•c)c equals??

OpenStudy (samigupta8):

@kainui

OpenStudy (samigupta8):

Actually i want to solve it without opening up a•b and a•c things!! And then solving for u,v and w individually and hence solve the three vectors and then find the volume of parallelepiped having these as its edges.. I want it to solve by another way out . Can anybody help pls??

OpenStudy (kindaljoshconall):

need help

OpenStudy (kindaljoshconall):

bud

OpenStudy (phi):

I assume you know the volume is given by the scalar triple product A dot (B cross C) which is also the "determinant" of A,B, and C made into a 3x3 matrix

OpenStudy (samigupta8):

Yep ! I know !

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (samigupta8):

@hartnn

OpenStudy (phi):

the only idea I have is to use det(AB)= det(A)*det(B) the volume of a parallelepiped with sides u, v, w is the scalar triple product u x v dot w or if we make a matrix M with u, v, and w being the rows, volume = det(M) as u,v,w are linear combinations of a,b,c, we can write M as M= [ a b c] M2 where [a b c] is a matrix with a,b,c being its columns, and M2 holds the coefficients thus |M| = | [ a b c] | * det(M2)

OpenStudy (phi):

I don't see enough structure in M2 to simplify it much, but maybe there is ? for this problem a*a= 1 and b*b=1 , so I would use that simplification

OpenStudy (samigupta8):

Yeah ! I have done like this bt when we solve for a•c, b•c and a•b v would get a very yuckkkk term and then it would be further multiplied with a,b and c so this is the worst part here!!

OpenStudy (samigupta8):

@phi

OpenStudy (samigupta8):

@vishweshshrimali5

OpenStudy (kindaljoshconall):

yea

OpenStudy (vishweshshrimali5):

The volume of the parallelepiped determined by the vectors a, b, and c is \(|c \cdot (a \times b)|\)

OpenStudy (samigupta8):

I know that bt i want to simplify the expression that r coming out to b very weird indeed! Can't we solve that in a very simplified manner.. When v would be just finding the dot product of a and b and b and c and a and c we would get terms to be very disgusting and not only this we would have to multiply them to a ,b and c once again to get u,v and w...

OpenStudy (phi):

I'm not sure what you are trying to do. Are you trying to find a formula for any 3 vectors a,b,c that then form u,v, and w ? btw, for this problem I'm getting \[ 13 \sqrt{3} -\frac{151}{8} \] which is not a very pretty number.